Renz enters NLEX at an average of 60 km/hr. Frederick enters NLEX 30 minutes later at an average of 80 km/hr. How long will it take Frederick to overtake Renz? How far will they be then from the moment they enter the expressway?
Answer by richwmiller(17219)

Gaining speed 80 -60= 20 kph
Headstart 60*0.5 =30 km.
Time to catch up 30/20=1.5 hours
Algebraic solution
60(t+0.5)=80t
60t+30.0=80t
30.0=80t-60t
30.0=20t
1.5=t
Distance to meeting 1.5*80=120 km.